OpenStudy (anonymous):

The domain of y = -2^x is X>2. True or False?

jimthompson5910 (jim_thompson5910):

can you plug in x = 0?

jimthompson5910 (jim_thompson5910):

and get a number as a result

OpenStudy (anonymous):

Therefore, it would be true? =)

jimthompson5910 (jim_thompson5910):

no, it's not true because x = 0 is definitely allowed This is because y = -2^x = -2^0 = -1 So x = 0 is definitely part of the domain

jimthompson5910 (jim_thompson5910):

So the domain is NOT x > 2

jimthompson5910 (jim_thompson5910):

Therefore, it's false

OpenStudy (anonymous):

Thank you so much! That makes sense now!

jimthompson5910 (jim_thompson5910):

you're welcome